A 0.065 M solution of the weak acid HA has a pH of 3.6 at 25°C .
(a) what is the ionization constant of the acid at 25°C, the percentage ionization of HA, and the ratio of A/AH in the solution.
(b) what is the pH of a 0.05 M HCL solution?
(c) How much stronger in acidity is the HCL solution than the HA solution?
(d) what is the Kb and pKb of the conjugate base, A?
1
Expert's answer
2019-10-21T07:17:04-0400

(a)

"[H^+] = [A^-]"

"[HA] = c(HA) - [H^+]"


"[H^+] = 10^{-pH} = 10^{-3.6} = 0.00025 M"

"k_a = {[H^+][A^-] \over [HA]} = {0.00025*0.00025 \over 0.065 - 0.00025} = 9.7*10^{-7}"

Ratio "A^-\/HA" in the solution:


"{[A^-] \over [HA]} = {0.00025 \over 0.065 - 0.00025} = 0.003861"

(b)

Chloric acid is a strong acid, the degree of dissociation is taken equal to 1, then:



"c(H^+) = c(HCl)"

"pH = -lg(c(H^+)) = -lg(0.05) = 1.3"

(c)

"{[H^+]_{HCl} \over [H^+]_{HA} } = {10^{-1.3} \over 10^{-3.6}} = 10^{2.3} = 199.5 ≈ 200"

(d)

The stronger the conjugate base, the weaker the acid.


"k_b = {k_w(H_2O) \over K_a(HA)} = {10^{-14} \over 9.7*10^{-7}} ≈ 10^6"

"pK_b = -lgK_b = -6"